: A force F = (30i – 40j) N is applied on an object, the object travels a displacement

S = (-9i-3j) m. Find the work done by the force on the object.
1
Expert's answer
2017-08-17T11:09:07-0400
W=F∙S = (30i-40j)∙(-9i-3j) = -270+120 = -150 J.